OpenStudy (anonymous):

Need help finding a scale factor betweeen simular shapes, going to be typing the problem below.

OpenStudy (anonymous):



Directrix (directrix):

Yes, 10 to 8 or 10/8 which equals 5/4. I usually express scale factor as a ratio. Of course, 5/4 = 1.25. Note that 5/4 = 10/8 = 5/4 which is the scale factor.